Assume the following sales data for a company: Current year Preceding year $757,611 $562,022 What is the percentage increase in sales from the preceding year to the current year? Select the correct answer. 074.18% 025.82% 0134.8096 034.80%
Based on the following data for the current year, what is the number of days sales in inventory? Net sales on account during year Cost of goods sold during year Accounts receivable, beginning of year Accounts receivable, end of year Inventory, beginning of year Inventory, end of year Do not round interim calculations. Round your final answer up to the nearest whole day Select the correct answer $442,358 166,384 49,656 49,781 30,662 39,889 O5 days O109 days O365 days 077 days
media%2F931%2F931f6820-2a33-4ba9-95f3-44
media%2Fbc2%2Fbc2268c5-b9b5-4b00-8d28-4f
Answer #1

Q1)

Answer: 4th option

Percentage increase = {(Current year – Previous year) / Previous year} × 100

                                    = {(757,611 – 562,022) / 562,022} × 100

                                    = (195,589 / 562,022) × 100

                                    = 34.80%

Q2)

Answer: 4th option

Inventory turnover ratio = CGS / Average inventory

                                        = 166,384 / {(30,662 + 39,889) / 2}

                                        = 166,384 / 35,275.5

                                        = 4.72 times

Required number of days = 365 days in a year / Inventory turnover ratio

                                         = 365 / 4.72

                                         = 77 days

Q3)

Answer: 2nd option

Quick assets = Accounts receivable + Cash + Marketable securities

                      = 35,000 + 25,000 + 36,000

                      = 96,000

Q4)

Answer: 1st option

Percentage increase = {(Current year – Previous year) / Previous year} × 100

                                    = {(1,025,000 – 820,000) / 820,000} × 100

                                    = (205,000 / 820,000) × 100

                                    = 25%